OpenStudy (juliebeans):

Solve the equation. If necessary, round your answer to the nearest tenth. 9x2 = 49 2.3, –2.3 5.4, –5.4 7, –7 21, –21

OpenStudy (juliebeans):

@sammixboo

OpenStudy (mathstudent55):

First, subtract 49 from both sides. What do you get?

OpenStudy (juliebeans):

erhm 1,5.4

OpenStudy (mathstudent55):

\(9x^2 = 49\) Start by subtracting 49 from both sides: \(9x^2 - 49 = 0\)

OpenStudy (juliebeans):

ok

OpenStudy (mathstudent55):

Now factor the left side. The left side is a difference of squares. Do you know how to factor a difference of squares? It follows this pattern: \(a^2 - b^2 = (a + b)(a - b) \)

OpenStudy (mathstudent55):

You have \(9x^2 - 49\) \(9x^2\) is the square of 3x. 49 is the square of 7. You can rewrite \(9x^2 - 49\) as \((3x)^2 - 7^2\) which clearly shows the two squares.

OpenStudy (mathstudent55):

Now you factor the left side using the pattern I showed above.

OpenStudy (juliebeans):

Can we combine the 9x2 and 3x(2) togther?

OpenStudy (juliebeans):

@mathstudent55

OpenStudy (mathstudent55):

No. I'll show you with a drawing to make it easier to see how a difference of squares factors.

OpenStudy (juliebeans):

ok

OpenStudy (mathstudent55):